OpenStudy (anonymous):

Which of the following are undefined terms? point segment plane angle @charlesbennett

OpenStudy (anonymous):

@mathmath333

OpenStudy (anonymous):

@morganchavez

OpenStudy (mathmath333):

they are i think point lines and planes

OpenStudy (anonymous):

In my notes i have point and plane as undefined. I hope that helps

OpenStudy (mathmath333):

angle can be infinite but it is all the angles are limited to\[0\leq \theta\leq 360\] while line,plane and point cannot be

OpenStudy (mathmath333):

note here that segment is not equal to line line and segment are two different terms